# Asymptotic stability of composite waves of two viscous shocks for relaxed compressible Navier-Stokes equations

### v2

@@ 4.1. $L^{2}$ energy estimates
$\leq C(\delta_{0}+\varepsilon_{1})\int_{0}^{1}|w_{1}|\mathrm{d}y_{1}+\frac{C}{\delta_{1}}\int_{\mathbb{R}}|(\widetilde{\Pi}_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x,$
@@
$\left(\Big{|}\frac{2M}{\sigma_{m}^{2}}\int_{0}^{1}w_{1}\mathrm{d}y_{1}\Big{|}-|\dot{X}_{1}|\right)^{2}\leq C(\delta_{0}+\varepsilon_{1})^{2}\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1}+\frac{C}{\delta_{1}^{2}}\left(\int_{\mathbb{R}}|(\widetilde{\Pi}_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x\right)^{2}.$
@@
$-\frac{\delta_{1}}{2M}|\dot{X}_{1}|^{2}\leq$ $-\frac{M\delta_{1}}{\sigma_{m}^{4}}\left(\int_{0}^{1}w_{1}\mathrm{d}y_{1}\right)^{2}+C\delta_{1}(\delta_{0}^{2}+\varepsilon_{1}^{2})\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1}.$
@@
$\leq C\delta_{1}\lambda_{1}\int_{\mathbb{R}}\frac{|p((\widetilde{v}_{1})^{X_{1}}_{x})|}{\delta_{1}}|(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x.$
@@
$C\delta_{1}\lambda_{1}\int_{\mathbb{R}}\frac{|p((\widetilde{v}_{1})^{X_{1}}_{x})|}{\delta_{1}}|(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x=C\delta_{1}\lambda_{1}\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1}.$
@@
$B_{11}-\mathcal{G}_{1}\leq C\delta_{1}\lambda_{1}\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1}.$
@@
Similarly, we only estimates the case of $i=1$. Using (2.11), (4.8), (4.10), (4.11) and Lemma 4.1, we have
$B_{21}$ $=\sigma_{1}\delta_{1}\int_{0}^{1}a(t,x+X_{1}(t))\frac{1}{p^{\prime}(\widetilde{v}_{1})}p(v|\widetilde{v})^{-X_{1}}\mathrm{d}y_{1}$
@@
On the other hand, define
$D(U):=\int_{\mathbb{R}}\frac{a}{\gamma p(v)}|\partial_{x}(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x.$
Using the methods in Lemma 4.5 of [25], we can get
$-D(U)\leq-\sum\limits_{i=1}\limits^{2}\delta_{i}\alpha_{m}(1-C(\delta_{0}+\varepsilon_{1}))\int_{0}^{1}y_{i}(1-y_{i})|\partial_{y_{i}}w_{i}|^{2}\mathrm{d}y_{i}.$
@@
$B_{1}+B_{2}-\mathcal{G}-\frac{3}{2}D$
$\leq\sum\limits_{i=1}\limits^{2}\delta_{i}\alpha_{m}\left((1+C(\delta_{0}+\varepsilon_{1}))\int_{0}^{1}|w_{i}|^{2}\mathrm{d}y_{i}-\frac{3}{4}(1-C_{0}(\delta_{0}+\varepsilon_{1}))\int_{0}^{1}y_{i}(1-y_{i})|\partial_{y_{i}}w_{i}|^{2}\mathrm{d}y_{i}\right),$
@@
and using the smallness of $\delta_{0},\delta_{1},\delta_{2},\varepsilon_{1}$, we derive that
@@
$\leq\sum\limits_{i=1}\limits^{2}\left(-\frac{\delta_{i}\alpha_{m}}{8}\int_{0}^{1}|w_{i}|^{2}\mathrm{d}y_{i}+\frac{5\delta_{i}\alpha_{m}}{4}\left(\int_{0}^{1}w_{i}\mathrm{d}y_{i}\right)^{2}\right).$
@@
$\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{2M}|\dot{X}_{i}|^{2}+B_{1}+B_{2}-G_{2}-\frac{3}{2}D$
@@
$\qquad\left.+\sigma_{i}\int_{\mathbb{R}}(a_{i})^{X_{i}}_{x}\frac{1+\gamma}{3\gamma}p(\widetilde{v})^{-\frac{1}{\gamma}-2}(p(v)-p(\widetilde{v}))^{3}\mathrm{d}x\right)$
@@
$\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{2M}|\dot{X}_{i}|^{2}+B_{1}+B_{2}-G_{2}-\frac{3}{2}D$
$\leq C\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}\left(-|(\widetilde{v}_{i})^{X_{i}}_{x}||(p(v)-p(\widetilde{v}))|^{2}+|(a_{i})^{X_{i}}_{x}||p(v)-p(\widetilde{v})|^{3}+|(a_{i})^{X_{i}}_{x}||\widetilde{v}-\widetilde{v}_{i}^{X_{i}}||p(v)-p(\widetilde{v})|^{2}\right)\mathrm{d}x.$
@@
$\leq C(-\mathcal{G}^{s}+K_{1}+K_{2})-\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{4M}|\dot{X}_{i}|^{2}+\sum\limits_{i=1}\limits^{2}\frac{C}{\delta_{i}}\sum\limits_{j=3}\limits^{8}|Y_{ij}|^{2}+\sum\limits_{i=3}\limits^{7}B_{i}-G_{1}-G_{3}-G+\frac{3}{2}D,$
@@
$\mathcal{G}^{s}:=\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x,\quad K_{1}:=\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}|(a_{i})^{X_{i}}_{x}||p(v)-p(\widetilde{v})|^{3}\mathrm{d}x,$
@@
For $K_{1}$, using the interpolation inequality, Hölder inequality, Young’s inequality, (2.13), Lemma 2.1 and $\lambda_{i}\leq C\sqrt{\delta_{i}}$, we have
$K_{1}$ $\leq\sum\limits_{i=1}\limits^{2}\frac{C\lambda_{i}}{\delta_{i}}\|p(v)-p(\widetilde{v})\|^{2}_{L^{\infty}}\sqrt{\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x}\sqrt{\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}|\mathrm{d}x}$
$\leq\sum\limits_{i=1}\limits^{2}\frac{C\lambda_{i}}{\sqrt{\delta_{i}}}\|\partial_{x}(p(v)-p(\widetilde{v}))\|_{L^{2}}\|p(v)-p(\widetilde{v})\|_{L^{2}}\sqrt{\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x}$
@@
$Y_{i4}\leq C\int_{\mathbb{R}}|\widetilde{v}-\widetilde{v}_{i}^{X_{i}}||(\widetilde{v}_{i})^{X_{i}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x$ $\leq C\delta_{0}\sqrt{\delta_{i}}\sqrt{\mathcal{G}^{s}}.$
@@
$\leq C\varepsilon_{1}^{2}\mathcal{G}^{s}.$
@@
$D$ $\leq C\int_{\mathbb{R}}|\partial_{x}(v-\widetilde{v})|^{2}\mathrm{d}x+C\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}|(\widetilde{v}_{i})_{x}^{X_{i}}|^{2}|p(v)-p(\widetilde{v})|^{2}\mathrm{d}x$
$\leq C_{2}\underbrace{\int_{\mathbb{R}}|\partial_{x}(v-\widetilde{v})|^{2}\mathrm{d}x}_{:=\mathcal{D}}++C(\delta_{1}^{2}+\delta_{2}^{2})\mathcal{G}^{s}.$
@@
$B_{4}\leq C\sum\limits_{i=1}\limits^{2}\left(\lambda_{i}\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}|^{\frac{1}{2}}|p(v)-p(\widetilde{v})||\Pi-\widetilde{\Pi}|\mathrm{d}x\right)\leq C(\lambda_{1}+\lambda_{2})\left(\mathcal{G}^{s}+G\right).$
@@
$\leq C(\delta_{1}+\delta_{2})\left(\mathcal{G}^{s}+G\right).$

### v3

@@ 4.1. $L^{2}$ energy estimates
$\leq C(\delta_{0}+\varepsilon_{1})\int_{0}^{1}|w_{1}|\mathrm{d}y_{1}+\frac{C}{\delta_{1}}\int_{\mathbb{R}}\Big(\phi_{2}|(\widetilde{v}_{1})^{X_{1}}_{x}|+\phi_{1}|(\widetilde{\Pi}_{1})^{X_{1}}_{x}|\Big)|p(v)-p(\widetilde{v})|\mathrm{d}x,$
@@
$\left(\Big|\frac{2M}{\sigma_{m}^{2}}\int_{0}^{1}w_{1}\mathrm{d}y_{1}\Big|-|\dot{X}_{1}|\right)^{2}\leq$ $C(\delta_{0}+\varepsilon_{1})^{2}\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1}+\frac{C}{\delta_{1}^{2}}\left(\int_{\mathbb{R}}\phi_{2}|(v_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x\right)^{2}$
$+\frac{C}{\delta_{1}^{2}}\left(\int_{\mathbb{R}}\phi_{1}|(\widetilde{\Pi}_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x\right)^{2}.$
@@
$\frac{C}{\delta_{1}^{2}}\left(\int_{\mathbb{R}}\phi_{2}|(v_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x\right)^{2}$
$\leq\frac{C}{\delta_{1}^{2}}\int_{\mathbb{R}}(\phi_{2}|(v_{1})^{X_{1}}_{x}|)^{2}\mathrm{d}x\int_{\mathbb{R}}|p(v)-p(\widetilde{v})|^{2}\mathrm{d}x$
$\leq\frac{C}{\delta_{1}^{2}}\|\phi_{2}|(v_{1})^{X_{1}}_{x}|\|_{L^{\infty}}\int_{\mathbb{R}}|(v_{1})^{X_{1}}_{x}|\mathrm{d}x\int_{\mathbb{R}}Q(v|\widetilde{v})\mathrm{d}x$
$\leq C\delta_{1}e^{-C\delta_{1}t}\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x.$
and
@@
$-\frac{\delta_{1}}{2M}|\dot{X}_{1}|^{2}\leq$ $-\frac{M\delta_{1}}{\sigma_{m}^{4}}\left(\int_{0}^{1}w_{1}\mathrm{d}y_{1}\right)^{2}+C\delta_{1}(\delta_{0}^{2}+\varepsilon_{1}^{2})\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1}+C\delta_{1}^{2}e^{-C\delta_{1}t}\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x.$
@@
$\leq C\delta_{1}\lambda_{1}\int_{\mathbb{R}}\frac{|p((\widetilde{v}_{1})^{X_{1}}_{x})|}{\delta_{1}}|\phi_{1}(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x+C\lambda_{1}\int_{\mathbb{R}}|(\widetilde{v}_{1})^{X_{1}}_{x})|\phi_{2}^{2}|p(v)-p(\widetilde{v})|^{2}\mathrm{d}x.$
@@
$C\delta_{1}\lambda_{1}\int_{\mathbb{R}}\frac{|p((\widetilde{v}_{1})^{X_{1}}_{x})|}{\delta_{1}}|\phi_{1}(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x=C\delta_{1}\lambda_{1}\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1}.$
On the other hand,
$C\lambda_{1}\int_{\mathbb{R}}|(\widetilde{v}_{1})^{X_{1}}_{x})|\phi_{2}^{2}|p(v)-p(\widetilde{v})|^{2}\mathrm{d}x\leq C\lambda_{1}\delta_{1}^{2}e^{-C\delta_{1}t}\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x.$
@@
$B_{11}-\mathcal{G}_{1}\leq C\delta_{1}\lambda_{1}\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1}+C\lambda_{1}\delta_{1}^{2}e^{-C\delta_{1}t}\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x.$
@@
Similarly, we only estimates the case of $i=1$. First, we have
$B_{21}=\sigma_{1}\int_{\mathbb{R}}a(\widetilde{v}_{1})^{X_{1}}_{x}\phi_{1}^{2}p(v|\widetilde{v})\mathrm{d}x+\sigma_{1}\int_{\mathbb{R}}a(\widetilde{v}_{1})^{X_{1}}_{x}(1-\phi_{1}^{2})p(v|\widetilde{v})\mathrm{d}x.$
Together with (2.13), (4.8), (4.10), (4.11) and Lemma 4.1, we get
$\sigma_{1}\delta_{1}\int_{0}^{1}a(t,x+X_{1}(t))\frac{\phi_{1}(x)^{2}}{p^{\prime}(\widetilde{v}_{1})}p(v|\widetilde{v})^{-X_{1}}\mathrm{d}y_{1}$
@@
and
$\sigma_{1}\int_{\mathbb{R}}a(\widetilde{v}_{1})^{X_{1}}_{x}(1-\phi_{1}^{2})p(v|\widetilde{v})\mathrm{d}x$ $=\sigma_{1}\int_{\mathbb{R}}a(\widetilde{v}_{1})^{X_{1}}_{x}(1+\phi_{1})\phi_{2}p(v|\widetilde{v})\mathrm{d}x$
$\leq C\int_{\mathbb{R}}|\widetilde{v}_{1})^{X_{1}}_{x}|\phi_{2}|p(v)-p(\widetilde{v})|^{2}\mathrm{d}x$
$\leq C\delta_{1}^{2}e^{-C\delta_{1}t}\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x.$
@@
For $D(U)$, noting that $\phi_{1}+\phi_{1}=1$, we derive that
$D(U)=\int_{\mathbb{R}}\frac{a\mu}{\gamma p(v)}(\phi_{1}+\phi_{2})|\partial_{x}(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x\geq\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}\frac{a\mu}{\gamma p(v)}\phi_{i}^{2}|\partial_{x}(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x.$
For any $\delta_{\ast}\in(0,1)$ small enough, Young’s inequality yields
$\int_{\mathbb{R}}\frac{a\mu}{\gamma p(v)}|\partial_{x}(\phi_{i}(p(v)-p(\widetilde{v})))|^{2}\mathrm{d}x$
$\leq(1+\delta_{\ast})\int_{\mathbb{R}}\frac{a\mu}{\gamma p(v)}\phi_{i}^{2}|\partial_{x}(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x+\frac{C}{\delta_{\ast}}\int_{\mathbb{R}}\frac{a\mu}{\gamma p(v)}|\partial_{x}\phi_{i}|^{2}|(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x,$
then, it holds
$-D(U)$ $\leq-\frac{1}{1+\delta_{\ast}}\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}\frac{a\mu}{\gamma p(v)}|\partial_{x}(\phi_{i}(p(v)-p(\widetilde{v})))|^{2}\mathrm{d}x+\frac{C}{\delta_{\ast}}\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}\frac{a\mu}{\gamma p(v)}|\partial_{x}\phi_{i}|^{2}|(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x$
$=:D_{1}+D_{2}.$
Following the similar method of Lemma 4.5 in [21], we derive
$D_{1}\leq-\sum\limits_{i=1}\limits^{2}\delta_{i}\alpha_{m}(1-C(\delta_{0}+\varepsilon_{1}))\int_{0}^{1}y_{i}(1-y_{i})|\partial_{y_{i}}w_{i}|^{2}\mathrm{d}y_{i}.$ (4.12)
From (3.28) and (3.8), the following holds for $D_{2}$ for any $(t,x)\in(0,T]\times\mathbb{R}$, it holds
$|\partial_{x}\phi(t,x)|\leq\frac{4}{\sigma_{2}-\sigma_{1}}\frac{1}{2},\quad i=1,2.$
Thus,
$D_{2}\leq\frac{C}{\delta_{\ast}t^{2}}\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x.$ (4.13)
Therefore, combining (4.12) and (4.13), we conclude that
$D(U)\leq-\sum\limits_{i=1}\limits^{2}\delta_{i}\alpha_{m}(1-C(\delta_{0}+\varepsilon_{1}))\int_{0}^{1}y_{i}(1-y_{i})|\partial_{y_{i}}w_{i}|^{2}\mathrm{d}y_{i}+\frac{C}{\delta_{\ast}t^{2}}\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x.$
@@
$B_{1}+B_{2}-\mathcal{G}-\frac{3}{4}D$
$\leq\sum\limits_{i=1}\limits^{2}\delta_{i}\alpha_{m}\left((1+C(\delta_{0}+\varepsilon_{1}))\int_{0}^{1}|w_{i}|^{2}\mathrm{d}y_{i}-\frac{3}{4}(1-C_{0}(\delta_{0}+\varepsilon_{1}))\int_{0}^{1}y_{i}(1-y_{i})|\partial_{y_{i}}w_{i}|^{2}\mathrm{d}y_{i}\right)$
$\quad+C\left(\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}e^{-C\delta_{i}t}+\frac{1}{\delta_{\ast}t^{2}}\right)\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x,$
@@
and using the smallness of $\delta_{0},\delta_{1},\delta_{2},\delta_{\ast},\varepsilon_{1}$, we derive that
@@
$\leq\sum\limits_{i=1}\limits^{2}\left(-\frac{\delta_{i}\alpha_{m}}{8}\int_{0}^{1}|w_{i}|^{2}\mathrm{d}y_{i}+\frac{5\delta_{i}\alpha_{m}}{4}\left(\int_{0}^{1}w_{i}\mathrm{d}y_{i}\right)^{2}\right)+C\left(\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}e^{-C\delta_{i}t}+\frac{1}{t^{2}}\right)\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x.$
@@
$\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{2M}|\dot{X}_{i}|^{2}+B_{1}+B_{2}-G_{2}-\frac{3}{4}D$
@@
$\qquad\left.+\sigma_{i}\int_{\mathbb{R}}(a_{i})^{X_{i}}_{x}\frac{1+\gamma}{3\gamma}p(\widetilde{v})^{-\frac{1}{\gamma}-2}(p(v)-p(\widetilde{v}))^{3}\mathrm{d}x\right)+C\left(\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}e^{-C\delta_{i}t}+\frac{1}{t^{2}}\right)\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x$
@@
$\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{2M}|\dot{X}_{i}|^{2}+B_{1}+B_{2}-G_{2}-\frac{3}{4}D$
$\leq C\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}\left(-|(\widetilde{v}_{i})^{X_{i}}_{x}||\phi_{i}(p(v)-p(\widetilde{v}))|^{2}+|(a_{i})^{X_{i}}_{x}||p(v)-p(\widetilde{v})|^{3}+|(a_{i})^{X_{i}}_{x}||\widetilde{v}-\widetilde{v}_{i}^{X_{i}}||p(v)-p(\widetilde{v})|^{2}\right)\mathrm{d}x$
$\quad+C\left(\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}e^{-C\delta_{i}t}+\frac{1}{t^{2}}\right)\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x.$
@@
$\leq C(-\mathcal{G}^{s}+K_{1}+K_{2})+C\left(\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}e^{-C\delta_{i}t}+\frac{1}{\delta_{\ast}t^{2}}\right)\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x$
$\quad-\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{4M}|\dot{X}_{i}|^{2}+\sum\limits_{i=1}\limits^{2}\frac{C}{\delta_{i}}\sum\limits_{j=3}\limits^{8}|Y_{ij}|^{2}+\sum\limits_{i=3}\limits^{7}B_{i}-G_{1}-G_{3}-G+\frac{3}{4}D,$
@@
$\mathcal{G}^{s}:=\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||\phi_{i}(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x,\quad K_{1}:=\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}|(a_{i})^{X_{i}}_{x}||p(v)-p(\widetilde{v})|^{3}\mathrm{d}x,$
@@
For $K_{1}$, firstly, we have
$\int_{\mathbb{R}}|(a_{i})^{X_{i}}_{x}||p(v)-p(\widetilde{v})|^{3}\mathrm{d}x\leq\frac{C\lambda_{i}}{\delta_{i}}\int_{\mathbb{R}}|(\widetilde{v})^{X_{i}}_{x}|(\phi_{i}+1-\phi_{i})|p(v)-p(\widetilde{v})|^{3}\mathrm{d}x,$
then, using the interpolation inequality, Hölder inequality, Young’s inequality, (2.15), Lemma 2.1 and $\lambda_{i}\leq C\sqrt{\delta_{i}}$, we get
$\frac{\lambda_{i}}{\delta_{i}}\int_{\mathbb{R}}|(\widetilde{v})^{X_{i}}_{x}|\phi_{i}|p(v)-p(\widetilde{v})|^{3}\mathrm{d}x$
$\leq\frac{C\lambda_{i}}{\delta_{i}}\|p(v)-p(\widetilde{v})\|^{2}_{L^{\infty}}\sqrt{\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||\phi_{i}(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x}\sqrt{\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}|\mathrm{d}x}$
$\leq\frac{C\lambda_{i}}{\sqrt{\delta_{i}}}\|\partial_{x}(p(v)-p(\widetilde{v}))\|_{L^{2}}\|p(v)-p(\widetilde{v})\|_{L^{2}}\sqrt{\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||\phi_{i}(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x}$
@@
On the other hand,
$\frac{\lambda_{i}}{\delta_{i}}\int_{\mathbb{R}}|(\widetilde{v})^{X_{i}}_{x}|(1-\phi_{i})|p(v)-p(\widetilde{v})|^{3}\mathrm{d}x\leq C\varepsilon_{1}\lambda_{i}\delta_{i}e^{-C\delta_{i}t}\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x.$
Thus, it holds
$K_{1}\leq C\varepsilon_{1}(D+\mathcal{G}^{s})+C\sum\limits_{i=1}\limits^{2}\varepsilon_{1}\lambda_{i}\delta_{i}e^{-C\delta_{i}t}\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x.$
@@
$Y_{i4}$ $\leq C\int_{\mathbb{R}}|\widetilde{v}-\widetilde{v}_{i}^{X_{i}}||(\widetilde{v}_{i})^{X_{i}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x$
$\leq C\sqrt{\int_{\mathbb{R}}(|\widetilde{v}-\widetilde{v}_{i}^{X_{i}}||(\widetilde{v}_{i})^{X_{i}}_{x}|)^{2}\mathrm{d}x}\sqrt{\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x}$
$\leq C\sqrt{\delta_{i}}\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}\sqrt{\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x}.$
@@
$\leq C\varepsilon_{1}^{2}\left(\int_{\mathbb{R}}|(\widetilde{v}_{i})_{x}^{X_{i}}||\phi_{i}(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x+\int_{\mathbb{R}}|(\widetilde{v}_{i})_{x}^{X_{i}}|(1-\phi_{i}^{2})|p(v)-p(\widetilde{v})|^{2}\mathrm{d}x\right)$
$\leq C\varepsilon_{1}^{2}\left(\mathcal{G}^{s}+\delta_{i}^{2}e^{-C\delta_{i}t}\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x\right).$
@@
$D$ $\leq(1+\kappa)\int_{\mathbb{R}}\frac{\mu(p^{\prime}(v))^{2}}{\gamma p(v)}|\partial_{x}(v-\widetilde{v})|^{2}\mathrm{d}x+C\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}|(\widetilde{v}_{i})_{x}^{X_{i}}|^{2}|p(v)-p(\widetilde{v})|^{2}\mathrm{d}x$
$\leq(1+\kappa)\int_{\mathbb{R}}\frac{\mu|p^{\prime}(v)|}{v_{m}-C\varepsilon_{1}}|\partial_{x}(v-\widetilde{v})|^{2}\mathrm{d}x+C\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}\left(\mathcal{G}^{s}+\delta_{i}^{2}e^{-C\delta_{i}t}\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x\right).$
@@
$B_{4}$ $\leq C\sum\limits_{i=1}\limits^{2}\left(\lambda_{i}\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}|^{\frac{1}{2}}|p(v)-p(\widetilde{v})||\Pi-\widetilde{\Pi}|\mathrm{d}x\right)$
$\leq C\sum\limits_{i=1}\limits^{2}\lambda_{i}\left(\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||p(v)-p(\widetilde{v})|^{2}\mathrm{d}x+G\right)$
$\leq C\sum\limits_{i=1}\limits^{2}\lambda_{i}\left(\mathcal{G}^{s}+\delta_{i}^{2}e^{-C\delta_{i}t}\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x+G\right).$
@@
$\leq C\sum\limits_{i=1}\limits^{2}\lambda_{i}\left(\mathcal{G}^{s}+\delta_{i}^{2}e^{-C\delta_{i}t}\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x+G\right).$